2. Sketch the path traced by r = 4 - 2t y = t - 9t and eliminate the parameters
Answer #1

= 4-2 2 얀 온 휘 TABLE t 기 - 3 -14 4 그이 - 2 - , 2 8 0 4 0 2. - - 8 2 4 -10 2 -14 0 GKAPH 개 (4,0) XV E19)Again a= 4-24² 2 y= ť- at 7 2+² = 4-1 t² = 4-1 2 ta (42) now put t value in y egh. So y = 4-2 2 2 914-11 2 Domain of above Gr